OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or equal to x - 5 y greater than or equal to -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (anonymous):

@amistre64 @Callisto @phi

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

Did you graph them and found the solution region?

OpenStudy (anonymous):

No, How would I do that?

OpenStudy (anonymous):

use graph paper graph the lines y=x-5 and y=-x-4

OpenStudy (aravindg):

yep

OpenStudy (anonymous):

find the region below or on the line y=x-5 and above or on the line y=-x-4

OpenStudy (aravindg):

you could also use a graphing calculator

OpenStudy (anonymous):

I'm in a library, so I don't have graphing paper, and I have no clue how to use a graphing calculator. :/

OpenStudy (aravindg):

its easy just plugin the quations in boxes y1 and y2

OpenStudy (aravindg):

*equations

OpenStudy (anonymous):

Okay. I'll try it right now;

OpenStudy (anonymous):

or go to https://www.desmos.com/calculator and type in the two equations (copy and paste works too)

OpenStudy (anonymous):

How do I make a less than or equal too sign on the graphing calculator?

OpenStudy (anonymous):

you don't need to, just graph the line and remember you are looking at the region less than or equal to it

OpenStudy (aravindg):

In my calculator click on the square having equal to after the y1 or y2

OpenStudy (anonymous):

I'm so confused > . <

OpenStudy (anonymous):

in desmos, enter the less than sign then the equals sign <= or >=

OpenStudy (anonymous):

Okay, I'm gonna use desmos then.

OpenStudy (aravindg):

ok.

OpenStudy (anonymous):

I did it. What next?

OpenStudy (anonymous):

you typed in both your equations?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

look at where your four points are

OpenStudy (anonymous):

see which one satisfies both inequalities

OpenStudy (anonymous):

0,5 and -4,5

OpenStudy (anonymous):

@Peter14

OpenStudy (amistre64):

i woulda just plugged in the 4 points to see what made true results

OpenStudy (amistre64):

(-5, 2) is 2 <= -5 - 5 true? is 2 >= --5 - 4 true? (5, -2) is -2 <= 5 - 5 true? is -2 >= -5 - 4 true? (-5, -2) is -2 <= -5 - 5 true? is -2 >= --5 - 4 true? (5, 2) is 2 <= 5 - 5 true? is 2 >= -5 - 4 true?

OpenStudy (anonymous):

Oh, okay! @amistre64 I'll try that,

OpenStudy (amistre64):

it is nice to know that other stuff, but it takes way to much time when taking a test.

OpenStudy (anonymous):

It's -5, -2, right?

OpenStudy (amistre64):

(-5, -2) -2 <= 0 true -2 >= 1 false not that one

OpenStudy (amistre64):

well, -10 up top, but still false :)

OpenStudy (anonymous):

So, it's -5, something right?

OpenStudy (amistre64):

dunno, lets try another one (5, -2) -2 <= 5 - 5 -2 <= 0 true -2 >= -5 - 4 -2 >= -9 true that fits nicely

OpenStudy (anonymous):

Awesome. Thanks :)